OpenStudy (yuii):

Given the function f(x) = 4 - 2x: If the domain is {-4, 0, 5}, find the range. A) {-4, 4, -6} B) {-4, 4, 14} C) {12, 4, -6} D) {12, 4, 14)

OpenStudy (yuii):

C?

OpenStudy (kittiwitti1):

f(x)=4-2x f(-4) = 4 - 2(-4) = 4 - (-8) = 12 f(0) = 4 - 2(0) = 4-0 = 4 f(5) = 4 - 2(5) = 4-10 = -6 Yup \(\checkmark\)

OpenStudy (kittiwitti1):

This is the last question I can help on, I have to go to class. Good luck on your studies!

OpenStudy (yuii):

Thank you for your help.

OpenStudy (shirleyarnett):

i agree with kittiwitti1 that the answer is c.

OpenStudy (yuii):

Thank you both.